OpenStudy (anonymous):

Find the first six terms of the sequence. (7 points) a1 = 1, an = 4 • an-1 (a)0, 4, 4, 8, 12, 16 (b)1, 4, 8, 12, 16, 20 (c)1, 4, 16, 64, 256, 1024 (d)4, 16, 64, 256, 1024, 4096

OpenStudy (anonymous):

simply replace the value of n into the sequence

OpenStudy (anonymous):

replace n=1 first

OpenStudy (anonymous):

Answer is C, I think. It is a recursive sequence. \[a _{n} = 4\times a _{n-1}\] and \[a _{1} = 1\] so \[a _{2} = a _{1}\times 4 = 4\] \[a _{3} = a _{2}\times 4 = 16\] \[a _{4} = a _{3}\times 4 = 64\] ans so on.